Starting in 7 years and 9 months you want to be able to withdraw $1700 at the beginning of every month. You deposit $100 000.00 immediately and then let it grow at a rate of 6.41% compounded quarterly. For how many years will you be able to withdraw these payments?


1
Expert's answer
2021-10-25T19:13:44-0400

To find compound interest we use:

"A=P(1+\frac{r}{n})^{nt}"

Principal"=" 100,000.00

r"=" "\frac{6.41}{400}=0.016025"

nt"=" 31 quarters

"A=100000(1+0.016025)^{31}"

"A=163,695.3037"

For Payout annuity we use the formula

"P_0= \frac{PMT(1-(1+\frac{r}{n})^{-nt}}{\frac{r}{n}}"

"P_0=" Account balance at the beginning(163,695.30)

PMT"=" Regular withdrawal amount(1700)

r"=" Annual interest rate(decimal)(0.0641)

n"=" Number of compounds per year(12)

t"=" Number of years to withdraw(?)


"163,695.30= \frac{1700(1-(1+\frac{0.0641}{12})^{-12t}}{\frac{0.0641}{12}}"

"\frac{163,695.30×0.0053}{1700}=1-1.0053^{-12t}"

"0.514356=1-1.0053^{-12t}"

"1.0053^{-12t}=1-0.514356"

"log 1.0053^{-12t}=log 0.485644"

"-12t=\frac{log0.485644}{log1.0053}"

"-12t=136.64"

"t= 11.38 years"